# Neuromorphic Intermediate Representation: A Unified Instruction Set for Interoperable Brain-Inspired Computing

### v1

@@ 4.4.1 LIF
where $\tau$ is the time constant, $v_{\text{leak}}$ is the membrane leak, $R$ is the resistance and $i(t)$ is the input current. Discretized using Euler’s method, this turns into:
$v(t+1)=\left(1-\frac{dt}{\tau}\right)v(t)+\frac{dt}{\tau}v_{\text{leak}}+\frac{dt}{\tau}Ri(t+1)$ (8)
@@
The course of the membrane potential $v$ can be exactly calculated for a constant input current $i=i_{0}=\text{const.}$. and initial condition $v(0)=v_{0}$ [46, eq. 1.7]:
@@
In Exponential-Euler mode, the decay factor $\alpha_{\text{decay}}$ is calculated such that in case without input ($i=0$ and $i_{\text{offset}}=0$) and with $v_{\mathrm{leak}}=0$ the voltage exactly matches the continuous-time dynamics:
@@
The LIF model in Rockpool is a time-discretised neuron membrane, with update equations
@@ 4.4.2 CuBa-LIF
In Forward-Euler mode we match the discrete Cuba-LIF equations (40-41) so that SpiNNaker2 parameters become:
@@
The bit-shift decay operation in Eqs. 60 and 61 approximates exponential decay for $V_{mem}$ and $I_{syn}$ state variables. When the integer quantised decay parameters $d_{*}$ obtained from $\tau_{*}$ are close to the corresponding floating-point values for $d_{*}$, this approximation causes only small deviations in dynamics. For values of $V_{mem}$ and $I_{syn}$ close to zero, where the bit-shift decay operation would not result in a decrease of the state variables, we include a unit decay per $dt$. We have neglected neuron indices in the above equations; all neurons have independent state variables and parameters.
4.5 Training setup
We proceed to describe the training setup used to obtain the architecture and parameters for the SCNN and SRNN in the second and third experiment, respectively. The training code can be found on https://github.com/neuromorphs/NIR/tree/paper/paper.
4.5.1 Spiking convolutional neural network
We trained a non-spiking ReLU based convolutional neural network on the neuromorphic MNIST dataset (N-MNIST) [47], that was later transferred to an equivalent spiking convolutional network with Sinabs, using the model transfer approach from  [48]. The last layer of the network contains 10 neurons, each one representing one digit, trained to have the largest firing rate for the neuron corresponding to the correct label. The training was done using backpropagation through time and surrogate gradients. Further details are available in our code repository.
4.5.2 Spiking recurrent neural network
Data from pressure readings encoded using a sigma-delta modulator with $\vartheta=1$ was used, accounting for a time binning with bin size equal to 5 ms. Compared to the original implementation, we introduced two simplifications to fit the connectivity and size constraints of the Xylo chip. First, by avoiding input copies and by collapsing the ON and OFF channels of each tactile sensor into a single spike array through an OR-like operation at every timestep, we reduced the input size to twelve. Second, we selected a subset of seven possible characters (’Space’, ’A’, ’E’, ’I’, ’O’, ’U’, ’Y’), thus making this a spatio-temporal classification problem with seven classes. For the recurrent layer of the spiking recurrent neural network (SRNN), 38 (zero, bias) or 40 (subtractive, no bias) CuBa-LIF neurons were used. To identify optimal hyperparameters for the SRNN, an optimization procedure adapted from the one described in [49] was performed. The network was trained with backpropagation-through-time using surrogate gradients in snnTorch. To optimize the spiking activity of the output neurons, training was performed on the cross entropy of the spike count at the output layer. The spike function was implemented with the Heaviside function in the forward pass and approximated with the fast sigmoid function in the backward pass [50]: $S\approx\frac{U}{1+k|U|}$ with $\frac{\partial S}{\partial U}=\frac{1}{(1+k|U|)^{2}}$ where $k=5$. A regularization term was added to this loss function to penalize both the $\ell^{1}$ norm of the average number of spikes per neuron and the $\ell^{2}$ norm of the total number of spikes in the recurrent layer. The coefficients for the regularization terms were chosen through the above-mentioned hyperparameter optimization, and the optimal values found were $\mu_{L1}=6\times 10^{-4}$, $\mu_{L2}=4\times 10^{-6}$ (zero, bias) and $\mu_{L1}=1\times 10^{-3}$, $\mu_{L2}=1\times 10^{-6}$ (subtractive, no bias). The objective was minimized using the Adam optimizer with a learning rate of 0.005 (zero, bias) or 0.001 (subtractive, no bias). The training proceeded for 500 epochs, and the parameters that yielded the highest validation accuracy were chosen.
5 Acknowledgements
The authors would like to acknowledge the Telluride Neuromorphic Cognition Engineering Workshop for providing a conducive environment for conceiving and implementing the first steps towards NIR. J.E.P. acknowledges funding from the EC Horizon 2020 Framework Programme under Grant Agreements 785907 and 945539 (HBP) as well as the Pioneer Centre for AI, under the Danish National Research Foundation grant number P1. S.A. acknowledges funding from the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Skłodowska-Curie Grant Agreement No. 860360 (POST DIGITAL). M.J. acknowledges funding by the German Research Foundation (DFG, Deutsche Forschungsgemeinschaft) as part of Germany’s Excellence Strategy – EXC 2050/1 – Project ID 390696704 – Cluster of Excellence ”Centre for Tactile Internet with Human-in-the-Loop” (CeTI) of Technische Universität Dresden. BV acknowledges funding by the German Federal Ministry for Economic Affairs and Climate Action (BMWK) under the contract 01MN23004F (ESCADE). V.F. acknowledges funding by the 3A-ITALY project (CUP E13C22001900001). G.U. acknowledges funding by the Ebrains-Italy project (CUP B51E22000150006). J.K.E. acknowledges funding from the National Science Foundation under Award ECCS-2332166. D.R.M and S.S acknowledge the ECSEL Joint Undertaking grant agreements 826655 “TEMPO” and 876925 “ANDANTE”; the KDT Joint Undertaking grant agreement 101097300 “EdgeAI”; Innosuisse and the Swiss State Secretariat for Education, Research and Innovation (SERI).
6 Contributions
J.E.P., S.A., M.J., F.C.B., B.V., G.L. and S.S. contributed to the NIR Core library. J.E.P., S.A., M.J., V.F., F.C.B., B.V., G.L., D.R.M., P.Z., K.H., T.C.S., J.K.E. and S.S. contributed to the conversions and experiments for the different library and hardware backends. J.E.P., S.A., B.V., S.S., G.U. and V.F. conceived the experiments. J.E.P., S.A., M.J., and J.K.E. contributed to the figures. All authors contributed to writing the paper.
7 Competing interests
The authors declare no competing interests.
8 Data availability
All data and code used in this paper is available at https://github.com/neuromorphs/nir.

### v2

@@ A.1 LIF
where $\tau$ is the time constant, $v_{\text{leak}}$ is the membrane leak, $R$ is the resistance and $i(t)$ is the input current. Discretized using explicit Euler’s method, this turns into:
$v(t+1)=\left(1-\frac{dt}{\tau}\right)v(t)+\frac{dt}{\tau}\,v_{\text{leak}}+\frac{dt}{\tau}R\ i(t)$ (9)
@@
The course of the membrane potential $v$ can be exactly calculated for a constant input current $i=i_{0}=\text{const.}$. and initial condition $v(0)=v_{0}$ [103, eq. 1.7]:
@@
In Exponential-Euler mode, the decay factor $\alpha_{\text{decay}}$ is calculated such that in case without input ($i=0$ and $i_{\text{offset}}=0$) and with $v_{\mathrm{leak}}=0$ the voltage exactly matches the analytical expression for continuous-time dynamics:
@@
The LIF model in Rockpool is a time-discretized neuron membrane, with difference or update equations
@@ A.2 CuBa-LIF
In Forward-Euler mode we match the discrete CuBa-LIF equations (41-42) so that SpiNNaker2 parameters become:
@@
The bit-shift decay operation in Eqs. 61 and 62 approximates exponential decay for $V_{mem}$ and $I_{syn}$ state variables. When the integer quantised decay parameters $d_{*}$ obtained from $\tau_{*}$ are close to the corresponding floating-point values for $d_{*}$, this approximation causes only small deviations in dynamics. For values of $V_{mem}$ and $I_{syn}$ close to zero, where the bit-shift decay operation would not result in a decrease of the state variables, we include a unit decay per $dt$. We have neglected neuron indices in the above equations; all neurons have independent state variables and parameters.
Appendix B Relation to other intermediate representation and compilation frameworks
To clarify the relation of NIR to other compilers, code generators, optimizers, and network modelling tools, we provide the following comparison table. Most of the frameworks in the table solve different problems. As such, the table provides high-level and independent descriptions of each framework to situate NIR in the literature.
Table 3: Comparison of ONNX, MLIR, LLVM, PyNN, NeuroML.
Feature / Framework ONNX (Open Neural Network Exchange) MLIR (Multi-Level Intermediate Representation) LLVM PyNN NeuroML NIR
Scope AI model formats and runtimes Compiler infrastructure targeting CPUs, GPUs, and more Compiler for the optimization and generation of machine code Abstraction layer for neural network descriptions, focusing on spiking models Models neural connectivity and dynamics, focusing on biological realism Declaration of discrete- or continuous-time neural primitives
Language Not applicable (format) Based on LLVM, supports many frontends (e.g., C++, Python) Primarily C++ Python XML-based description Not applicable (format)
Supported Platforms / Frameworks TensorFlow, PyTorch, and other major AI frameworks Integrates with TensorFlow, PyTorch, and others through LLVM Wide range of hardware from desktops to embedded devices NEURON, Brian, NEST NEURON, Brian, jLEMS Lava, Loihi, Nengo, Norse, Rockpool, Sinabs, snnTorch, SpiNNaker2, Speck, Spyx, Xylo
Typical Use Cases Exchanging models between different AI frameworks Optimizing performance of machine learning models at compile time Low-level optimization of applications in various programming languages Research and educational use in computational neuroscience Detailed simulations of neuronal behavior and neural circuitry for research Application-dependent models and also able to exchange models between platforms